If a carpenter wants to be sure that the corner of a building is square and measures out 6 ft and 8 ft along the sides, how long should she make the diagonal?

Answer by laoman(51)

The diagonal would make the angle a right angle, therefore we can use pythagoras theorem to solve this one
6^2+8^2= d^2
36+64 = d^d
d^2 = 100
d
=10
She should make the diagonals 10 ft